the perimeter of a rectangle is 48. The width is 2 more than half the length. find the dimensions

Answer by algebrahouse.com(1659)

x = length
(x/2) + 2 = width {width is 2 more than half the length}

Perimeter of a rectangle is 2(width) + 2(length}

2x + 2[(x/2) + 2] = 48 {substituted width and length into perimeter formula}
2x + x + 4 = 48 {used distributive property}
3x + 4 = 48 {combined like terms}
3x = 44 {subtracted 4 from each side}
x = 44/3 {divided each side by 3}
(x/2) + 2 = 28/3 {substituted 44/3, in for x, into (x/2) + 2}

length is 44/3
width is 28/3

For more help from me, visit: www.algebrahouse.com